OpenStudy (anonymous):

@jacob194485 ok one more Factor completely −5x^3 + 10x^2. A. Prime B. −x(5x2 − 10x) C. 5x(−x2 + 2x) D. −5x2(x − 2) but this time dont give me the answer i need to learn how to do it!_! plz!

OpenStudy (anonymous):

i am tagging every one on my fans list lol

OpenStudy (anonymous):

@jordan123321 @Micah1mcdugle @roz_mar23 @MinightTheWolfie @alanwallace

OpenStudy (anonymous):

that r online

OpenStudy (bee_see):

what do they have in common?

OpenStudy (bee_see):

factor that out

OpenStudy (anonymous):

the only thing i see is the x

OpenStudy (anonymous):

(-5) is common to both coefficients x^2 is common to both variables (- 5x^2) (x - 2)

OpenStudy (anonymous):

r u on the question or the answer choices

OpenStudy (anonymous):

the questions. that hasjust been explained

OpenStudy (anonymous):

now, to factor out, you must look at the two or more independent variables

OpenStudy (anonymous):

you must check for the co-efficient and also the powers

OpenStudy (anonymous):

THE X'S ?

OpenStudy (anonymous):

yes, x is the independent variable separated by + sign. now note before u can factor out, there must be a separation with either + or -

OpenStudy (anonymous):

R THE INDEPENDENT VARIABLES THE X'S?

OpenStudy (anonymous):

ohh ohkay

OpenStudy (anonymous):

the variable involved is just X

OpenStudy (anonymous):

now, lets look at some examples before working on that,..... facor out .. \[x ^{4}+x\]

OpenStudy (anonymous):

can you try factorize that?

OpenStudy (anonymous):

ok i will try

OpenStudy (anonymous):

x(x^4)??

OpenStudy (anonymous):

hmmmm.wrong

OpenStudy (anonymous):

now have you heard of distribution property?

OpenStudy (anonymous):

distribution property states that a(b+c)=ab +ac

OpenStudy (anonymous):

its just a way of opening the bracket

OpenStudy (anonymous):

ummm ive hered tha but i thought it was a+b=c^2

OpenStudy (anonymous):

nooooooo

OpenStudy (anonymous):

it says that 2(1+3) is same as 2(1)+2(3)

OpenStudy (anonymous):

i am sooo soorry for being so dumb and frustrating

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

try that and tell me if u get thesame thing

OpenStudy (anonymous):

yes is that the foil presedure

OpenStudy (anonymous):

just a start to it

OpenStudy (anonymous):

now, look at \[x ^{4} +x\]

OpenStudy (anonymous):

ok i am looking lol

OpenStudy (anonymous):

first, they are asking if there is a common thing between x^4 and x

OpenStudy (anonymous):

the x ??

OpenStudy (anonymous):

now recal that the things to watch out are 1) the variables and power, @ 2) the co-efficient of the two separated variables

OpenStudy (anonymous):

now, what is the co-efficient of x^4?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

do you have idea?

OpenStudy (anonymous):

sorry had to do something

OpenStudy (anonymous):

can you answer the question ?

OpenStudy (anonymous):

what is the co-efficient of x^4?

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

i am so confused i will give you medal any way

OpenStudy (anonymous):

please stay let me teach you and i promise you will understand. never give up

OpenStudy (anonymous):

@i<3u43v4

OpenStudy (anonymous):

i am triing but it is hard i hat f*cking school ok go on

OpenStudy (anonymous):

the co-efficient of a variable is the number beside or multiplying that veriable

OpenStudy (anonymous):

5x. the co-effiecient of that 5x or x is the 5

OpenStudy (anonymous):

so it is 4

OpenStudy (anonymous):

here the co-effiecient of \[x ^{4}\]

OpenStudy (anonymous):

is 1

OpenStudy (anonymous):

there is no number 1 tho

OpenStudy (anonymous):

note this, every variable or number is multiplied by invisible 1 and devided by invisible 1

OpenStudy (anonymous):

e.g \[1x/1\]

OpenStudy (anonymous):

ok so it is always 1 right?

OpenStudy (anonymous):

but in maths is better to say x instead of 1x/1

OpenStudy (anonymous):

i have a question for you. \[2\times1=\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2

OpenStudy (anonymous):

any thing times 1 is it self

OpenStudy (anonymous):

now what that means is that 2 is times d by 1 and at the same time divided by 1, this is how it ought to be \[(2\times1)\div1\]

OpenStudy (anonymous):

but in maths we will just take 2 rather than\[ (2\times 1)\div1\]

OpenStudy (anonymous):

now did you understand that step?

OpenStudy (anonymous):

did you understand why we always say 2 instead of (2X1)/1?

OpenStudy (anonymous):

waw. later den

OpenStudy (anonymous):

yes i under stand now!!

OpenStudy (anonymous):

it is d?

OpenStudy (anonymous):

@GIL.ojei

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

thanks to openstudy.

OpenStudy (anonymous):

Sorry i never got the notification if you still need the answer its −5x2(x − 2)